OpenStudy (anonymous):

Which of the following is the expansion of (x + 11)^4? A) x^4 + 4x^3 + 6x^2 + 4x + 1 B) None of the Above C) 1 + 44x + 726x^2 + 5324x^3 + 14641x^4 D) x^4 + 11x^3 + 121x^2 + 1331x + 14641

OpenStudy (anonymous):

I get none of the above? X^4 + 44x^3 + 726 x^2 + 5324 X + 13641

OpenStudy (anonymous):

thank you. any input, nikvist?

OpenStudy (asnaseer):

the first term in the expansion must be x^4 and the last term must be 11^4=121^2=14641 so you can say from this that either B) or D) is the answer do you know how to use the binomial expansion?

OpenStudy (asnaseer):

this might help you: http://www.mathsrev.com/binomial-expansion/